Solve the initial value problem
y"(t)+ 2y'(t)+y(t)=3e^(-t) ; y'(0)=2 y(0)=4
1
Expert's answer
2020-05-21T16:47:57-0400

"y''+2y'+y=3e^{-t}\\
y(0)=4, y'(0)=2\\"

Solution of equation find in form

"y=y_1+y_2"

Consider equation

"y''+2y'+y=0"

Write characteristic equation

"\lambda^2+2\lambda+1=0 \to \lambda_1=\lambda_2=-1"

then the solution of equation is

"y_1=c_1e^{-t}+c_2te^{-t}"

The solution "y_2" find in form

"y_2=at^2e^{-t}"

input "y_2" in the first equation

"2ae^{-t}-2tae^{-t} -2tae^{-t} +t^2ae^{-t}+\\
+4tae^{-t}-2t^2ae^{-t}+t^2ae^{-t}=3e^{-t}\\
2a=3 , a=1.5"

so

"y_2=1.5t^2e^{-t}"

and

"y=c_1e^{-t}+c_2te^{-t}+1.5t^2e^{-t}"

Find

"y'=-c_1e^{-t}+c_2e^{-t}-c_2te^{-t}+\\
+2t\cdot 1.5e^{-t}-1.5t^2e^{-t}\\
y(0)=c_1 \to c_1=4\\
y'(0)=-c_1+c_2 \to -c_1+c_2=2 \to\\
-4+c_2=2 \to c_2=6"

So

"y=4e^{-t}+6te^{-t}+1.5t^2e^{-t}"